# Equivariant $\mathcal{D}$-stability for Actions of Tensor Categories

### v1

@@ 4. A $\mathcal{C}$-equivariant McDuff type result
For a Hilbert $A$-bimodule $E$ and a Hilbert $B$-bimodule $H$ we denote their external tensor product by $E\otimes H$ which is a Hilbert $A\otimes B$-bimodule (see [27]), where $A\otimes B$ denotes the minimal tensor product of the $*{}^{*}$-algebras $A$ and $B$.
@@
Let $(G,J):\mathcal{C}\curvearrowright A$ be an action and $B$ be a $*{}^{*}$-algebra. We denote by $G\otimes\mathrm{id}_{B}:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A\otimes B)$ the $\mathbb{C}$-linear functor defined by
@@
Following [39, Definition 1.3], a unital $*{}^{*}$-algebra $\mathcal{D}\not\cong\mathbb{C}$ is strongly self-absorbing if there is an isomorphism $\mathcal{D}\cong\mathcal{D}\otimes\mathcal{D}$ that is approximately unitary equivalent to the first factor embedding. In light of Remark 4.2 it makes sense to introduce the following generalisation of [42, Definition 1.7].
@@
Let $(G,J):\mathcal{C}\curvearrowright A$ be an action and $\mathcal{D}$ be a strongly self-absorbing $*{}^{*}$-algebra. We say $(G,J)$ is equivariantly $\mathcal{D}$-stable if $(G,J)$ is cocycle conjugate to $(G\otimes\mathrm{id}_{\mathcal{D}},J\otimes 1_{\mathcal{D}})$, i.e. there exists a cocycle conjugacy between $(G,J)$ and $(G\otimes\mathrm{id}_{\mathcal{D}},J\otimes 1_{\mathcal{D}})$.
@@
Let $A$ be a separable $*{}^{*}$-algebra, $\mathcal{C}$ be a unitary tensor category with countably many isomorphism classes of simple objects and $(G,J):\mathcal{C}\curvearrowright A$ be an action. Let $\mathcal{D}$ be a strongly self-absorbing $*{}^{*}$-algebra. The following are equivalent:
@@
    $(A,G,J)$ is cocycle conjugate to $(A,G\otimes\mathrm{id}_{\mathcal{D}^{\otimes\infty}},J\otimes 1_{\mathcal{D}^{\otimes\infty}})$;
@@
by $\varphi(d\otimes d^{\prime})=\psi(d)\otimes d^{\prime}$. [^1] We view $\psi(d)\otimes d^{\prime}$ as an element of the codomain by lifting $\psi(d)$ to an element in $\bigcap_{X\in\mathrm{Irr}(\mathcal{C})}A\cap G(X)^{\prime}$, tensoring on $d^{\prime}$, and then descending to the quotient. It is easy to see that the result is independent of the choice of lift. For the remainder of the proof write
(4.6) $B=\frac{\bigcap_{X\in\mathrm{Irr}(\mathcal{C})}(A\otimes\mathcal{D})_{\infty}\cap(G(X)\otimes 1_{\mathcal{D}})^{\prime}}{(A\otimes\mathcal{D})_{\infty}\cap(A\otimes 1_{\mathcal{D}})^{\perp}},$
which is a $*{}^{*}$-subalgebra of $F(A\otimes 1_{\mathcal{D}},(A\otimes\mathcal{D})_{\infty})$. [^2] This follows in the same way as Proposition 3.2. Firstly, note that for any $b\in F(A\otimes 1_{\mathcal{D}},(A\otimes\mathcal{D})_{\infty})$ and $\xi\in G(X)\otimes\mathcal{D}$ with $X\in\mathrm{Irr}(\mathcal{C})$ the actions $b\rhd\xi$ and $\xi\lhd b$, contained in $(G(X)\otimes\mathcal{D})_{\infty}$, are well defined as $G(X)$ is non-degenerate. Moreover, for $d\in\mathcal{D}$ and $\xi\in G(X)$ one has that $\varphi(1_{\mathcal{D}}\otimes d)\rhd\xi\otimes 1_{\mathcal{D}}=(\psi(1_{\mathcal{D}})\otimes d)\rhd(\xi\otimes 1_{\mathcal{D}})=\xi\otimes d$ as $\psi$ is unital and $G(X)$ is non-degenerate. Similarly, $\varphi(d\otimes 1_{\mathcal{D}})\rhd(\xi\otimes 1_{\mathcal{D}})=(\psi(d)\otimes 1_{\mathcal{D}})\rhd(\xi\otimes 1_{\mathcal{D}})\subset G(X)_{\infty}\otimes 1_{\mathcal{D}}\subset(G(X)\otimes\mathcal{D})_{\infty}$.
@@
(4.7) $\|v^{*}(1_{\mathcal{D}}\otimes d)v-d\otimes 1_{\mathcal{D}}\|<\varepsilon$
for any $d\in\mathcal{F}_{1}$. Set $u=\varphi(v)$. Let $d\in\mathcal{F}_{1}$, $X\in\mathrm{Irr}(\mathcal{C})$ and $\xi\in G(X)$ with $\|\xi\|\leq 1$. Since $(\xi\otimes 1_{\mathcal{D}})\lhd\varphi(x)=\varphi(x)\rhd(\xi\otimes 1_{\mathcal{D}})$ for all $x\in\mathcal{D}\otimes\mathcal{D}$, we have
(4.8) $\begin{split}\mathrm{dist}(&G(X)_{\infty}\otimes 1_{\mathcal{D}},u^{*}\rhd(\xi\otimes d)\lhd u)\\ &=\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(v^{*})\rhd(\xi\otimes d)\lhd\varphi(v))\\ &=\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(v^{*}(1_{\mathcal{D}}\otimes d))\rhd(\xi\otimes 1_{\mathcal{D}})\lhd\varphi(v))\\ &=\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(v^{*}(1_{\mathcal{D}}\otimes d))\varphi(v)\rhd(\xi\otimes 1_{\mathcal{D}})\\ &=\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(v^{*}(1_{\mathcal{D}}\otimes d)v)\rhd(\xi\otimes 1_{\mathcal{D}}))\\ &<\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(d\otimes 1_{\mathcal{D}})\rhd(\xi\otimes 1_{\mathcal{D}}))+\varepsilon\\ &=\varepsilon,\end{split}$
where we used that $\varphi(d\otimes 1_{\mathcal{D}})\rhd(\xi\otimes 1_{\mathcal{D}})\in G(X)_{\infty}\otimes 1_{\mathcal{D}}$.
@@
  1. (1)
    $\|w_{n}\rhd(\xi\otimes 1_{\mathcal{D}})-(\xi\otimes 1_{\mathcal{D}})\lhd w_{n}\|\leq\varepsilon$ for all $\xi\in\mathcal{G}_{X},X\in K$,
  2. (2)
    $\mathrm{dist}(G(X)\otimes 1_{\mathcal{D}},w_{n}^{*}\rhd(\xi\otimes d)\lhd w_{n})\leq\varepsilon$, for all $\xi\in\mathcal{G}_{X}$, $d\in\mathcal{F}_{1}$ and $X\in K$.
Therefore, the cocycle morphism consisting of the $*{}^{*}$-homomorphism $\mathrm{id}_{A}\otimes 1_{\mathcal{D}}$ with family of linear maps $h^{X}=\mathrm{id}_{G(X)}\otimes 1_{\mathcal{D}}$ for $X\in\mathcal{C}$ satisfies the conditions of [13, Theorem 6.2] and (iv) follows. ∎
@@ 5. Application to $\mathcal{D}$-stability for stationary AF-actions
If $\mathcal{C}$ is a unitary fusion category, then an AF-action is a action on an AF-algebra which is an inductive limit of actions on finite dimensional $*{}^{*}$-algebras. These can be described by choosing, for each $n\in\mathbbm{N}$, a unitary finitely semisimple (left) $\mathcal{C}$-module category $\mathcal{M}_{n}$, a generating object $m_{n}\in\mathcal{M}_{n}$, and a $\mathcal{C}$-module functor $F_{n}:\mathcal{M}_{n}\rightarrow\mathcal{M}_{n+1}$ such that $F_{n}(m_{n})\cong m_{n+1}$. Note that $m_{n}$ is determined up to isomorphism by the choices of $\mathcal{M}_{n},F_{n}$ and $m_{0}$. This data is called an enriched Bratelli diagram for an AF-action and completely determines an AF-action up to equivalence (though it is not unique).
@@
We now unpack the definition of the inductive limit action from the enriched Bratteli diagram. We set $A_{n}:=\text{End}_{\mathcal{M}}(m_{0}\triangleleft Y^{\otimes n})$ for each $n\in\mathbb{N}$, which are finite dimensional $*{}^{*}$-algebras.
@@
Now, before we state our main result of this section, for any (strict) unitary tensor category $\mathcal{E}$ and any object $Y\in\mathcal{E}$, we define the AF-algebra $B^{Y}$ by setting $B^{Y}_{n}:=\text{End}_{\mathcal{E}}(Y^{\otimes n})$ with connecting maps $x\mapsto x\otimes 1_{Y}$, and taking the inductive limit.
@@
Let $\mathcal{C}$ be a unitary fusion category and $\mathcal{M}$ a unitary indecomposable $\mathcal{C}$-module category with generating object $m_{0}\in\mathcal{M}$. Let $Y\in\mathcal{C}^{*}_{\mathcal{M}}$, and consider the stationary AF action $G:\mathcal{C}\rightarrow\mathrm{Corr}(A)$ defined by $(m_{0},\mathcal{M},Y)$. Then $(A_{\infty}\cap A^{\prime})^{G}$ contains $B^{Y}$ as a unital subalgebra. In particular, if $B^{Y}$ contains a strongly self-absorbing unital subalgebra $\mathcal{D}$, then the action $G$ is $\mathcal{D}$-stable.
@@
Let $x\in B^{Y}_{m}=\text{End}_{\mathcal{C}^{*}_{\mathcal{M}}}(Y^{\otimes m})$. Define $x_{n}:=1_{m_{0}\triangleleft Y^{\otimes n}}\triangleleft x\in\text{End}(m_{0}\triangleleft Y^{\otimes(n+m)}):=A_{n+m}$. By construction, $x_{n}$ commutes with the image of $G_{n}(X)$ under $j_{n}^{X}\circ\cdots\circ j^{X}_{n+m-1}$. Hence, the the map $x\mapsto(x_{n})_{n=1}^{\infty}$ induces a unital injective $*{}^{*}$-homomorphism $B^{Y}_{m}\rightarrow(A_{\infty}\cap A^{\prime})^{G}$ for each $m\in\mathbb{N}$. As these maps are compatible with the connecting maps $x\mapsto x\otimes 1_{Y}$, we obtain a unital injective $*{}^{*}$-homomorphism $B^{Y}\rightarrow(A_{\infty}\cap A^{\prime})^{G}$. ∎
We call a stationary AF-action satisfying the hypotheses of the above theorem an indecomposable stationary AF-action. Recall an object $Y$ in a fusion category $\mathcal{E}$ is called a strong tensor generator of $\mathcal{E}$ if there exists an $n$ such that every simple object of $\mathcal{E}$ is isomorphic to a summand of $Y^{\otimes n}$.
@@
Let $G$ be the indecomposable stationary AF-action of $\mathcal{C}$ on the unital AF-algebra $A$ determined by the triple $(m_{0},\mathcal{M},Y)$. If $Y$ is a strong tensor generator for $\mathcal{C}^{*}_{\mathcal{M}}$, then $G$ is $\mathcal{Z}$-stable.
@@
Since $Y$ is a strong tensor generator, the natural Bratteli diagram of $B^{Y}$ given by tensoring $Y^{\otimes n}$ has a strongly connected adjaceny matrix. It follows that $B^{Y}$ is a simple AF-algebra (see for example [6, Chapter 6]). Hence, $B^{Y}$ is $\mathcal{Z}$-stable. The result now follows by combining Theorem 5.2 with Theorem 4.4. ∎

### v2

@@ 4. A $\mathcal{C}$-equivariant McDuff type result
For a Hilbert $A$-bimodule $E$ and a Hilbert $B$-bimodule $H$ we denote their external tensor product by $E\otimes H$ which is a Hilbert $A\otimes B$-bimodule (see [27]), where $A\otimes B$ denotes the minimal tensor product of the C-algebras $A$ and $B$.
@@
Let $(G,J):\mathcal{C}\curvearrowright A$ be an action and $B$ be a C-algebra. We denote by $G\otimes\mathrm{id}_{B}:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A\otimes B)$ the $\mathbb{C}$-linear functor defined by
@@
Following [39, Definition 1.3], a unital C-algebra $\mathcal{D}\not\cong\mathbb{C}$ is strongly self-absorbing if there is an isomorphism $\mathcal{D}\cong\mathcal{D}\otimes\mathcal{D}$ that is approximately unitary equivalent to the first factor embedding. In light of Remark 4.2 it makes sense to introduce the following generalisation of [42, Definition 1.7].
@@
Let $(G,J):\mathcal{C}\curvearrowright A$ be an action and $\mathcal{D}$ be a strongly self-absorbing C-algebra. We say $(G,J)$ is equivariantly $\mathcal{D}$-stable if $(G,J)$ is cocycle conjugate to $(G\otimes\mathrm{id}_{\mathcal{D}},J\otimes 1_{\mathcal{D}})$, i.e. there exists a cocycle conjugacy between $(G,J)$ and $(G\otimes\mathrm{id}_{\mathcal{D}},J\otimes 1_{\mathcal{D}})$.
@@
Let $A$ be a separable C-algebra, $\mathcal{C}$ be a unitary tensor category with countably many isomorphism classes of simple objects and $(G,J):\mathcal{C}\curvearrowright A$ be an action. Let $\mathcal{D}$ be a strongly self-absorbing C-algebra. The following are equivalent:
@@
    $(A,G,J)$ is cocycle conjugate to $(A\otimes\mathcal{D}^{\otimes\infty},G\otimes\mathrm{id}_{\mathcal{D}^{\otimes\infty}},J\otimes 1_{\mathcal{D}^{\otimes\infty}})$;
@@
$(\mathrm{id}_{A}\otimes\psi,\{\mathrm{id}_{G(X)}\otimes\psi\}_{X\in\mathcal{C}})\circ(\varphi,\{h^{X}\}_{X\in\mathcal{C}})$
is the desired cocycle conjugacy.
@@
Let $(h_{n})_{n\in\mathbb{N}}$ be a sequential approximate unit for $A$ and denote by $\rho_{n}:\mathcal{D}\rightarrow\mathcal{D}^{\otimes\infty}$ the unital -homomorphism taking $\mathcal{D}$ to the $n$-th tensor entry of $\mathcal{D}^{\otimes\infty}$ i.e.  $\rho_{n}(d)$ is the image of $1_{\mathcal{D}}\otimes\ldots\otimes 1_{\mathcal{D}}\otimes d\in\mathcal{D}^{\otimes n}$ under the canonical inclusion $\mathcal{D}^{\otimes n}\subset\mathcal{D}^{\otimes\infty}$. Now, the mapping
$\phi:\mathcal{D}$ $\rightarrow F(A\otimes\mathcal{D}^{\otimes\infty})$
$d$ $\mapsto(h_{n}\otimes\rho_{n}(d))$
is a unital embedding into $F(A\otimes\mathcal{D}^{\otimes\infty})^{G\otimes\mathrm{id}_{\mathcal{D}^{\otimes\infty}}}$, as required.
@@
by $\varphi(d\otimes d^{\prime})=\psi(d)\otimes d^{\prime}$. [^1] We view $\psi(d)\otimes d^{\prime}$ as an element of the codomain by lifting $\psi(d)$ to an element in $\bigcap_{X\in\mathrm{Irr}(\mathcal{C})}A\cap G(X)^{\prime}$, tensoring on $d^{\prime}$, and then descending to the quotient. It is easy to see that the result is independent of the choice of lift. For the remainder of the proof we denote the codomain C-algebra of $\varphi$ by $B$. That $B$ is indeed a C-algebra follows in the same way as Proposition 3.2, note also that it is a subalgebra of $F(A\otimes 1_{\mathcal{D}},(A\otimes\mathcal{D})_{\infty})$.
Firstly, for any $b\in F(A\otimes 1_{\mathcal{D}},(A\otimes\mathcal{D})_{\infty})$ and $\xi\in G(X)\otimes\mathcal{D}$ with $X\in\mathrm{Irr}(\mathcal{C})$ we denote by $b\rhd\xi$ and $\xi\lhd b$ the elements of $(G(X)\otimes\mathcal{D})_{\infty}$ defined by lifting $b$ to an element of $(A\otimes\mathcal{D})_{\infty}\cap(A\otimes 1_{\mathcal{D}})^{\prime}$ and applying the left and right action by this lift to $\xi$ respectively. That this is independent of the lift follows as $G(X)$ is non-degenerate. Moreover, for $d\in\mathcal{D}$ and $\xi\in G(X)$ one has that
(4.6) $\varphi(1_{\mathcal{D}}\otimes d)\rhd\xi\otimes 1_{\mathcal{D}}=(\psi(1_{\mathcal{D}})\otimes d)\rhd(\xi\otimes 1_{\mathcal{D}})=\xi\otimes d$
as $\psi$ is unital and $G(X)$ is non-degenerate. Similarly,
(4.7) $\varphi(d\otimes 1_{\mathcal{D}})\rhd(\xi\otimes 1_{\mathcal{D}})=(\psi(d)\otimes 1_{\mathcal{D}})\rhd(\xi\otimes 1_{\mathcal{D}})\subset G(X)_{\infty}\otimes 1_{\mathcal{D}}.$
@@
(4.8) $\|v^{*}(1_{\mathcal{D}}\otimes d)v-d\otimes 1_{\mathcal{D}}\|<\varepsilon,\forall d\in\mathcal{F}_{1}.$
Set $u=\varphi(v)$. Let $d\in\mathcal{F}_{1}$, $X\in\mathrm{Irr}(\mathcal{C})$ and $\xi\in G(X)$ with $\|\xi\|\leq 1$. Combining (4.6), (4.7) and that $(\xi\otimes 1_{\mathcal{D}})\lhd\varphi(x)=\varphi(x)\rhd(\xi\otimes 1_{\mathcal{D}})$ for all $x\in\mathcal{D}\otimes\mathcal{D}$, we have
(4.9) $\begin{split}\mathrm{dist}(&G(X)_{\infty}\otimes 1_{\mathcal{D}},u^{*}\rhd(\xi\otimes d)\lhd u)\\ &=\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(v^{*})\rhd(\xi\otimes d)\lhd\varphi(v))\\ &=\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(v^{*}(1_{\mathcal{D}}\otimes d))\rhd(\xi\otimes 1_{\mathcal{D}})\lhd\varphi(v))\\ &=\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(v^{*}(1_{\mathcal{D}}\otimes d)v)\rhd(\xi\otimes 1_{\mathcal{D}}))\\ &<\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},\varphi(d\otimes 1_{\mathcal{D}})\rhd(\xi\otimes 1_{\mathcal{D}}))+\varepsilon\\ &=\varepsilon,\end{split}$
@@
Since $u$ is homotopic to to the identity, we may lift $u$ to a unitary $w$ in $(M(A)\otimes\mathcal{D})_{\infty}$. As $w$ lifts $u$, it follows that
$w\rhd(\xi\otimes 1_{\mathcal{D}})=(\xi\otimes 1_{\mathcal{D}})\lhd w$
and
$\mathrm{dist}(G(X)_{\infty}\otimes 1_{\mathcal{D}},w^{*}\rhd(\xi\otimes d)\lhd w)<\varepsilon$
@@
(4.11) $\|w_{n}$ $\rhd(\xi\otimes 1_{\mathcal{D}})-(\xi\otimes 1_{\mathcal{D}})\lhd w_{n}\|\leq\varepsilon,\ \text{for all}\ \xi\in\mathcal{G}_{X},X\in K,$
(4.12) $\mathrm{dist}$ $(G(X)\otimes 1_{\mathcal{D}},w_{n}^{*}\rhd(\xi\otimes d)\lhd w_{n})\leq\varepsilon,\ \text{for all}\ \xi\in\mathcal{G}_{X},$
$d\in\mathcal{F}_{1},\text{and}\ X\in K.$
Therefore, the cocycle morphism consisting of the -homomorphism $\mathrm{id}_{A}\otimes 1_{\mathcal{D}}$ with family of linear maps $h^{X}=\mathrm{id}_{G(X)}\otimes 1_{\mathcal{D}}$ for $X\in\mathcal{C}$ satisfies the conditions of [13, Theorem 6.2] and (iv) follows. ∎
@@ 5. Application to $\mathcal{D}$-stability for stationary AF-actions
If $\mathcal{C}$ is a unitary fusion category, then an AF-action is a action on an AF-algebra which is an inductive limit of actions on finite dimensional C-algebras. These can be described by choosing, for each $n\in\mathbbm{N}$, a unitary finitely semisimple (left) $\mathcal{C}$-module category $\mathcal{M}_{n}$, a generating object $m_{n}\in\mathcal{M}_{n}$, and a $\mathcal{C}$-module functor $F_{n}:\mathcal{M}_{n}\rightarrow\mathcal{M}_{n+1}$ such that $F_{n}(m_{n})\cong m_{n+1}$. Note that $m_{n}$ is determined up to isomorphism by the choices of $\mathcal{M}_{n},F_{n}$ and $m_{0}$. This data is called an enriched Bratelli diagram for an AF-action and completely determines an AF-action up to equivalence (though it is not unique).
@@
We now unpack the definition of the inductive limit action from the enriched Bratteli diagram. We set $A_{n}:=\text{End}_{\mathcal{M}}(m_{0}\triangleleft Y^{\otimes n})$ for each $n\in\mathbb{N}$, which are finite dimensional C-algebras.
@@
Now, before we state our main result of this section, for any (strict) unitary tensor category $\mathcal{E}$ and any object $Y\in\mathcal{E}$, we define the AF-algebra $B^{Y}$ as the inductive limit of $B^{Y}_{n}:=\text{End}_{\mathcal{E}}(Y^{\otimes n})$ with connecting maps $x\mapsto x\otimes 1_{Y}$.
@@
Let $\mathcal{C}$ be a unitary fusion category and $\mathcal{M}$ a unitary indecomposable $\mathcal{C}$-module category with generating object $m_{0}\in\mathcal{M}$. Let $Y\in\mathcal{C}^{*}_{\mathcal{M}}$, and consider the stationary AF action $G:\mathcal{C}\rightarrow\mathrm{Corr}(A)$ defined by $(m_{0},\mathcal{M},Y)$. Then $(A_{\infty}\cap A^{\prime})^{G}$ contains $B^{Y}=\lim\operatorname{End}_{\mathcal{C}^{*}_{\mathcal{M}}}(Y^{\otimes n})$ as a unital subalgebra. In particular, if $B^{Y}$ contains a strongly self-absorbing unital subalgebra $\mathcal{D}$, then the action $G$ is $\mathcal{D}$-stable.
@@
Let $x\in B^{Y}_{m}=\text{End}_{\mathcal{C}^{*}_{\mathcal{M}}}(Y^{\otimes m})$. Define $x_{n}:=1_{m_{0}\triangleleft Y^{\otimes n}}\triangleleft x\in\text{End}(m_{0}\triangleleft Y^{\otimes(n+m)}):=A_{n+m}$. By construction, $x_{n}$ commutes with the image of $G_{n}(X)$ under $j_{n}^{X}\circ\cdots\circ j^{X}_{n+m-1}$. Hence, the the map $x\mapsto(x_{n})_{n=1}^{\infty}$ induces a unital injective -homomorphism $B^{Y}_{m}\rightarrow(A_{\infty}\cap A^{\prime})^{G}$ for each $m\in\mathbb{N}$. As these maps are compatible with the connecting maps $x\mapsto x\otimes 1_{Y}$, we obtain a unital injective -homomorphism $B^{Y}\rightarrow(A_{\infty}\cap A^{\prime})^{G}$. ∎
We call a stationary AF-action satisfying the hypotheses of the above theorem an indecomposable stationary AF-action. Recall an object $Y$ in a fusion category $\mathcal{E}$ is called a strong tensor generator of $\mathcal{E}$ if there exists an $n\in\mathbb{N}$ such that every simple object of $\mathcal{E}$ is isomorphic to a summand of $Y^{\otimes n}$.
@@
Let $G$ be the indecomposable stationary AF-action of the unitary fusion category $\mathcal{C}$ on the unital AF-algebra $A$ determined by the triple $(m_{0},\mathcal{M},Y)$. If $Y$ is a strong tensor generator for $\mathcal{C}^{*}_{\mathcal{M}}$, then $G$ is $\mathcal{Z}$-stable.
@@
Since $Y$ is a strong tensor generator, there exists an $n\in\mathbb{N}$ such that every simple object of $\mathcal{C}^{*}_{\mathcal{M}}$ is isomorphic to a summand of $Y^{\otimes n}$. Therefore, the Bratteli diagram for $B^{Y}$ given by tensoring $Y^{\otimes n}$ at each stage has an adjacency matrix with all entries non-zero. It follows that $B^{Y}$ is a simple AF-algebra (see for example [6, Chapter 6]). Hence, $B^{Y}$ is $\mathcal{Z}$-stable. The result now follows by combining Theorem 5.2 with Theorem 4.4. ∎